Your great Aunt Matilda put some money in an account for you on the day you were born. This account pays 8% interest per year. On your 21st birthday the account balance was $5033.83. How much money did your great aunt Matilda originally put in the account?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

5033.83=P*(1.08)^21

P=5033.83/(1.08)^21

=$1000(Approx).